OpenStudy (toxicsugar22):

calculate the value o f(H30+) from the given (-0H) and label the solution as acidic or basic a. 10^-8 (H30+)=_________*10^M

TheSmartOne (thesmartone):

@dan815 You good in Chemistry?

OpenStudy (anonymous):

"Assuming that the solution is aqueous, you look for the following: 1)pH <7, acidic; pH>7, basic, pH=7, neutral. If you are give pOH, then pOH>7 is acidic and pOH<7 is basic; you can calculate pH from pOH by subtracting pOH from 14. If you are given concentrations in 10^-x form, simply remember that pH (for H3O+ concentrations) and pOH (for OH- concentrations) are simply the 10 base logarithm of the negative of the concentration (Ex. pH of [H3O+] = 10^-4 = 4). Therefore, 1) is basic, 2) is neutral, 3) is basic, 4) is acidic, 5) is acidic (you can have a pOH of 14 - and a corresponding pH of ~0), and 6) is acidic."

TheSmartOne (thesmartone):

@abb0t @dan815

OpenStudy (abb0t):

You know pH = -log[H\(^+\)]. And 14 = pH + pOH To solve for concentration, you eliminate the log, by raising both sides to the power of 10. This is basic algebra.

OpenStudy (anonymous):

Base from what abbot said. Known: [OH-] = 10^-8 Unknown: [H+] = ? This is like the 3 steps we did before, but instead of finding [OH], we need to find [H+] 1. \(\sf pOH= -log [OH-]\) -> use this to find pOH 2. Using the calculated pOH, find pH by this formula: \(\sf pH=14 -pOH\) 3. After finding pH, apply this: \([H^+]= 10^{-pH}\) same steps like we did before (:

OpenStudy (toxicsugar22):

is it -log( 10^-8)

OpenStudy (toxicsugar22):

to start the problem

OpenStudy (anonymous):

yes, just continue doing it... i'll be checking it

OpenStudy (anonymous):

p.s. don't delete your comments @toxicsugar22

OpenStudy (toxicsugar22):

14-8=6

OpenStudy (toxicsugar22):

is poH 8

OpenStudy (anonymous):

yes...

OpenStudy (toxicsugar22):

PH IS 6

OpenStudy (toxicsugar22):

RIGHT

OpenStudy (anonymous):

yep, then..?

OpenStudy (toxicsugar22):

{h+}=10^-6

OpenStudy (anonymous):

right (: now look at the \(\sf pH\), is it acidic or basic?

OpenStudy (toxicsugar22):

it is acidic

OpenStudy (anonymous):

yes! now , can you do the 2nd one now?

OpenStudy (toxicsugar22):

so how should i write the answer. should i say 1*10^-6

OpenStudy (anonymous):

yes, it's the same

OpenStudy (toxicsugar22):

thanks

OpenStudy (anonymous):

np

OpenStudy (toxicsugar22):

so the next one is 6.00*10^-9 M h30+=_____*10^ M

OpenStudy (toxicsugar22):

I got the anwer i think, I want you to check it

OpenStudy (toxicsugar22):

is it -log(6.00*10^-9 )

OpenStudy (toxicsugar22):

then POH is 8.221

OpenStudy (toxicsugar22):

am i correct so far?

OpenStudy (anonymous):

yes (:

OpenStudy (toxicsugar22):

then PH IS 5.779

OpenStudy (anonymous):

mhmmm.. keep going

OpenStudy (toxicsugar22):

THEN IS THE ANWER H+=10^-5779

OpenStudy (toxicsugar22):

WHICH IS 5.779

OpenStudy (toxicsugar22):

AND IT IS ACIDIC

OpenStudy (anonymous):

are you sure? \(\sf [H^+]= 10^{-5.779}=? \) I got a different answer

OpenStudy (anonymous):

try calculating it again (:

OpenStudy (toxicsugar22):

I GOT 5.779

OpenStudy (anonymous):

for [H+] ? if it's for pH, then it is 5.779

OpenStudy (toxicsugar22):

oh so ph is 5.779

OpenStudy (anonymous):

yes, then do redo step 3

OpenStudy (toxicsugar22):

so then h+ = 8.221

OpenStudy (toxicsugar22):

right

OpenStudy (toxicsugar22):

and its basic

OpenStudy (anonymous):

no no and no... let's recap step 1: you calculated pOH which is equal to 8.221 step 2: calculating pH which is 14-8.221 = 5.779 now do step 3: calculating \(\sf [H+]= 10^ {-5.779}\)

OpenStudy (toxicsugar22):

im getting 5.779

OpenStudy (toxicsugar22):

are you getting that

OpenStudy (toxicsugar22):

i put -log(10^-5.779) right

OpenStudy (anonymous):

no, no need to get the -log 'coz you eliminated it \(\sf pH= -log [H^+]\) evaluate this to solve for [H+] by raising both sides by base 10 therefore you'll have \(\sf [H^+]= 10^{-pH}\)

OpenStudy (anonymous):

if you calculated the log, you'll get the same result.

OpenStudy (toxicsugar22):

so it is 5.779

OpenStudy (toxicsugar22):

or no

OpenStudy (toxicsugar22):

or is it 1.6663

OpenStudy (anonymous):

not quite right, it is \(\sf 1 \times 10^{-5.779}\) or \(\sf 1.663 \times 10^{-6}\) calculate this

OpenStudy (anonymous):

lol no need to calculate, i already did :3 just try doing it and see if you get the same

OpenStudy (anonymous):

are you getting it?